# Multilingual DistilWhisper: Efficient Distillation of Multi-Task Speech Models via Language-Specific Experts

### v3

@@ 5 Results
DistilWhisper versus other adaptation approaches. For whisper-small, we observe that both FT and LoRA-FT approaches (middle portion of Table 1) are able to improve performance on both in- and out-of-domain test sets. However, for FT this boost in performance comes with the cost of language specialization. In contrast to that, LoRA-FT is a light adaptation technique that does not modify the pre-trained representation. This method increases performance on both in-domain (avg -13.1) and out-of-domain (avg -3.5) test sets compared to whisper-small. DistilWhisper further improves performance over whisper-small (avg -15.3) and LoRA-FT (avg -2.2) for in-domain data. It also presents better out-of-domain adaptation capabilities compared to LoRA-FT (avg -2.1).
@@
Impact of knowledge distillation. We observe that DistilWhisper on average outperforms all other adaptation approaches (FT, LoRA-FT) for in- and out-of-domain test sets (bottom portion of Table 1). Comparing our models (CLSR-FT and DistilWhisper), we observe that the version with KD (DistilWhisper) exhibits a slight increase in average in-domain performance (-0.3). In out-of-domain settings, this model consistently outperforms CLSR-FT across all languages (avg -0.6), which confirms our initial hypothesis that the KD loss leverages the robustness from the teacher into the final model. Overall, these results highlight the effectiveness of our proposed architecture: we were able to reduce the out-of-domain performance gap between whisper-large-v2 and whisper-small by (avg -5.5) with a parameter overhead at inference time of only (25 M).
@@
Effect of training data size. We now show the effectiveness of our approach on lower and higher data resource settings. For this, we select a subset of languages for which we find more training data available on CV-13 (ca, th, ta). Table 2 presents results for our approach in low (3k utterances; 4 h), and higher-resource settings (28k utterances; 40 h), compared to the 10k results from Table 1. We observe that, as expected, increasing the amount of trainable examples leads to superior ASR performance for both approaches, with the leveraging of KD (DistilWhisper) being consistently superior to CLSR-FT. For the 28k setup (ca, th, ta), we are able to reduce the out-of-domain WER gap between whisper-large-v2 and whisper-small by (from to WER).555whisper-large-v2 and whisper-small avg FLEURS scores for ca, th, ta are respectively 12.5 and 24.5. For the 3k setup, we reduce the WER gap by 35.8% using only 4 h of training data. This implies that our approach has the potential to improve ASR performance across low-resource languages for which less training data is available.
@@
Gate Activation Analysis. To better understand how the model uses routing mechanism, we plot gate activation statistics for both CLSR-FT and DistilWhisper on Figure 2. We observe that the models tend to rely more on the new LS modules in out-of-domain settings (FLEURS vs CV-13), which could be attributed to the greater complexity and larger size of sentences in FLEURS. Also, as expected, increasing the training data size leads to more reliable LS modules, and therefore higher LS usage. The only exception for this is Thai at the 28k setup, and this might be due to dataset quality and requires further investigation. When comparing the 3 languages, we observe that Catalan exhibits a higher reliance on LS routes, which could also be related to the data quality for this language in CV-13. Finally, we observe that for languages with a weaker teacher (Thai, Tamil) the model may receive contradictory signals at lower-resource settings (3k, 10k), leading to less LS routing usage with KD. However, in the higher resource setting (28k), KD usage leads systematically to more reliable LS module and therefore higher LS routing.
@@ 6 Conclusion
We presented DistilWhisper, a parameter-efficient distillation approach that boosts performance of whisper-small by leveraging the robustness from the whisper-large-v2 into a smaller model, while preserving its multilingual capabilities. This is done by adding language-specific gated modules, and by jointly optimizing ASR fine-tuning and KD losses. Compared to LoRA adapters, and across eight languages, we are able to consistently improve performance in both in- and out-of-domain test sets, while adding only a negligible number of parameters at inference time. We believe that such architecture would make usage of Whisper models accessible to a larger amount of researchers and practitioners since it allows to boost the performance of a low-inference cost model by 35.2% using only 14 h of training data.
@@ Acknowledgements
This work was partially funded by the ANR project Diké (grant number ANR-21-CE23-0026-02). It was also partially funded by the European Horizon 2022 project UTTER (Unified Transcription and Translation for Extended Reality), under grant agreement No 101070631.

### v4

@@ 5 Results
For whisper-small, we observe that both FT and LoRA-FT approaches (middle portion of Table 1) are able to improve performance on both in- and out-of-domain test sets. However, FT achieves this improvement at the cost of language specialization, reducing performance in other languages. In contrast to that, LoRA-FT is a light adaptation technique that does not modify the pre-trained representation. This method increases performance on both in-domain (avg -13.1) and out-of-domain (avg -3.5) test sets compared to whisper-small. DistilWhisper further improves performance over whisper-small (avg -15.3) and LoRA-FT (avg -2.2) for in-domain data. It also presents better out-of-domain adaptation capabilities compared to LoRA-FT (avg -2.1).
5.2 Impact of knowledge distillation
@@
We observe that DistilWhisper on average outperforms all other adaptation approaches (FT, LoRA-FT) for in- and out-of-domain test sets (bottom portion of Table 1). Comparing our models (CLSR-FT and DistilWhisper), we observe that the version with KD (DistilWhisper) exhibits a slight increase in average in-domain performance (-0.3). In out-of-domain settings, this model consistently outperforms CLSR-FT across all languages (avg -0.6), which confirms our initial hypothesis that the KD loss leverages the robustness from the teacher into the final model. Overall, these results highlight the effectiveness of our proposed architecture: we reduce the out-of-domain performance gap between whisper-large-v2 and whisper-small by (avg -5.5) with a parameter overhead at inference time of only (25 M).
5.3 Effect of training data size
@@
We now show the effectiveness of our approach on lower and higher data resource settings. For this, we select a subset of languages for which we find more training data available on CV-13 (ca, th, ta). Table 2 presents results for our approach in low (3k utterances; 4 h), and higher-resource settings (28k utterances; 40 h), compared to the 10k results from Table 1. We observe that, as expected, increasing the amount of trainable examples leads to superior ASR performance for both approaches, with the leveraging of KD (DistilWhisper) being consistently superior to CLSR-FT. For the 28k setup (ca, th, ta), we reduce the out-of-domain WER gap between whisper-large-v2 and whisper-small by (from to WER).555whisper-large-v2 and whisper-small avg FLEURS scores for ca, th, ta are respectively 12.5 and 24.5. For the 3k setup, we reduce the WER gap by 35.8% using only 4 h of training data. This implies that our approach has the potential to improve ASR performance across low-resource languages for which less training data is available.
5.4 Gate Activation Analysis
@@
To better understand how the model uses routing mechanism, we plot gate activation statistics for both CLSR-FT and DistilWhisper in Figure 1. We observe that the models tend to rely more on the new language-specific modules in out-of-domain settings (FLEURS vs CV-13), which could be attributed to the greater complexity and larger size of sentences in FLEURS. Also, as expected, increasing the training data size leads to more reliable LS modules, and therefore higher LS usage. The only exception for this is Thai at the 28k setup, and this might be due to dataset quality and requires further investigation. When comparing the 3 languages, we observe that Catalan exhibits a higher reliance on LS routes, which could also be related to the data quality for this language in CV-13. Finally, we observe that for languages with a weaker teacher (Thai, Tamil) the model may receive contradictory signals at lower-resource settings (3k, 10k), leading to less LS routing usage with KD. However, in the higher resource setting (28k), KD usage leads systematically to more reliable LS module and therefore higher LS routing.
@@ 6 Conclusion
We presented DistilWhisper, a parameter-efficient distillation approach that boosts performance of whisper-small by leveraging the robustness from the whisper-large-v2 into a smaller model, while preserving its multilingual capabilities. This is done by adding language-specific gated modules, and by jointly optimizing ASR fine-tuning and KD losses. Compared to LoRA adapters, and across eight languages, we are able to consistently improve performance in both in- and out-of-domain test sets, while adding only a negligible number of parameters at inference time. We believe this architecture makes Whisper models more accessible to researchers and practitioners, as it boosts the performance of a low-inference cost model by 35.2% using only 14 h of training data.
@@ Acknowledgements
This research was partially supported by the French Agence Nationale de la Recherche, ANR as part of the project Diké - Bias, Fairness and Ethics of Compressed Language Models, under grant number ANR-21-CE23-0026-02.
This work was also co-funded by the European Union’s Horizon Europe Research and Innovation programme through the project UTTER – Unified Transcription and Translation for Extended Reality666For more information please visit: https://he-utter.eu/. under Grant Agreement No. 101070631.
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Effect of Temperature and Distillation Loss
We explore two temperature settings (1 and 3), and compare Jensen-Shannon (JS) loss with the traditional Kullback-Leibler (KL) in the 28k setup. Results are presented for validation (Table 3) and test sets (Table 4).
| FLEURS | CV-13 | FLEURS | CV-13 | |||||
| avg | avg | ca | ta | th | ca | ta | th | |
| JS w/ | 14.8 | 8.8 | 12.7 | 10.7 | 12.9 | 5.9 | 18.7 | 9.9 |
| JS w/ | 15.4 | 8.5 | 14.5 | 10.0 | 12.7 | 6.0 | 18.9 | 9.4 |
| KL w/ | 15.6 | 10.2 | 15.1 | 13.8 | 12.8 | 6.3 | 18.8 | 10.5 |
| KL w/ | 15.7 | 8.6 | 15.2 | 10.5 | 13.0 | 5.9 | 18.8 | 9.5 |
| FLEURS | CV-13 | FLEURS | CV-13 | |||||
| avg | avg | ca | ta | th | ca | ta | th | |
| JS w/ | 15.4 | 9.3 | 13.1 | 19.2 | 14.0 | 11.3 | 10.9 | 5.7 |
| JS w/ | 16.3 | 9.7 | 14.8 | 20.1 | 14.1 | 11.8 | 11.3 | 5.9 |
| KL w/ | 15.6 | 10.8 | 14.6 | 18.7 | 13.3 | 14.9 | 11.3 | 6.2 |
| KL w/ | 16.5 | 9.7 | 15.8 | 19.8 | 14.0 | 12.2 | 11.1 | 5.9 |
We observe overall better validation scores using JS, a trend that is confirmed by our test results. Regarding temperature, JS equipped with presents inferior validation scores, but superior CV-13 (in-domain) test scores. In this work, we selected our models based on their performance on the validation set of CV-13, which is why we report results for JS with only.
Appendix B Stability of the Methods
| Domain | Split | CLSR-FT | JS () | KL () | |||||||||
| s1 | s2 | s3 | avg | s1 | s2 | s3 | avg | s1 | s2 | s3 | avg | ||
| Catalan (ca) | |||||||||||||
| FLEURS | Val | 13.1 | 13.9 | 13.6 | 13.50.4 | 12.7 | 12.6 | 13.1 | 12.80.3 | 15.1 | 14.8 | 14.7 | 14.80.2 |
| Test | 13.5 | 13.3 | 13.6 | 13.50.2 | 13.1 | 13.6 | 13.3 | 13.30.2 | 14.6 | 14.6 | 14.3 | 14.50.2 | |
| CV-13 | Val | 9.6 | 9.6 | 9.0 | 9.40.4 | 10.7 | 9.3 | 8.8 | 9.61.0 | 13.8 | 13.8 | 13.7 | 13.80.1 |
| Test | 11.4 | 11.3 | 11.2 | 11.30.1 | 11.3 | 11.3 | 11.3 | 11.30.03 | 14.9 | 14.6 | 15.0 | 14.80.2 | |
| Tamil (ta) | |||||||||||||
| FLEURS | Val | 19.2 | 18.9 | 19.0 | 19.10.2 | 18.7 | 19.1 | 18.5 | 18.80.3 | 18.1 | 17.7 | 17.7 | 17.80.2 |
| Test | 19.6 | 19.8 | 19.9 | 19.80.1 | 19.2 | 19.2 | 19.6 | 19.30.2 | 18.5 | 18.3 | 17.9 | 18.20.3 | |
| CV-13 | Val | 9.3 | 9.3 | 9.0 | 9.20.2 | 9.9 | 9.3 | 8.9 | 9.40.5 | 10.5 | 10.4 | 11.4 | 10.80.5 |
| Test | 11.3 | 11.4 | 11.2 | 11.30.1 | 10.9 | 10.9 | 11.2 | 11.00.2 | 11.4 | 11.2 | 9.8 | 10.80.9 | |
| Thai (th) | |||||||||||||
| FLEURS | Val | 13.8 | 13.6 | 13.5 | 13.60.2 | 12.9 | 12.9 | 12.9 | 12.90.02 | 12.8 | 13.0 | 13.0 | 12.90.1 |
| Test | 13.8 | 13.8 | 14.2 | 13.90.2 | 14.0 | 13.6 | 13.6 | 13.70.2 | 13.3 | 13.6 | 13.6 | 13.50.1 | |
| CV-13 | Val | 6.4 | 6.0 | 5.7 | 6.00.3 | 5.9 | 5.8 | 5.9 | 5.90.1 | 6.3 | 6.5 | 6.3 | 6.40.1 |
| Test | 6.2 | 6.0 | 5.9 | 6.00.2 | 5.7 | 5.8 | 5.8 | 5.70.1 | 6.2 | 6.3 | 6.2 | 6.20.1 | |
| Averages | |||||||||||||
| FLEURS | Val | 15.4 | 15.5 | 15.4 | 15.40.04 | 14.8 | 14.9 | 14.8 | 14.80.1 | 15.3 | 15.1 | 15.1 | 15.20.1 |
| Test | 15.6 | 15.6 | 15.9 | 15.70.2 | 15.4 | 15.5 | 15.5 | 15.50.03 | 15.5 | 15.5 | 15.2 | 15.40.2 | |
| CV-13 | Val | 8.4 | 8.3 | 7.9 | 8.20.3 | 8.8 | 8.1 | 7.9 | 8.30.5 | 10.2 | 10.2 | 10.5 | 10.30.2 |
| Test | 9.7 | 9.6 | 9.4 | 9.50.1 | 9.3 | 9.4 | 9.4 | 9.30.1 | 10.8 | 10.7 | 10.3 | 10.60.3 | |
We investigate robustness by repeating each approach — fine-tuned only (CLSR-FT), DistilWhisper with JS (), and DistilWhisper with KL ()—over three random seeds in the 28k setup. Table 5 reports per-language results and dataset averages as meanstd WER (↓) across seeds.
With a 95% confidence intervals on the macro-averaged scores across seeds, DistilWhisper with JS is statistically superior to the fine-tuning-only baseline, while KL is comparable to JS in mean but exhibits higher variability on some splits, with is probably related to the borderline effects reported in recent research (js_vs_kd_acl23; js_vs_kd_bb2023).77795% confidence intervals over seed means: CLSR-FT [15.54, 15.87], JS [15.42, 15.49], KL [15.24, 15.57]. In short, JS () offers the best trade-off between accuracy and stability across languages and domains, confirming that distillation improves over CLSR-FT with lower seed sensitivity.